OpenStudy (anonymous):

Write an equation in point-slope form of the line that passes through (5, 0), and whose slope is -1/2.

OpenStudy (anonymous):

A. y = -1/2(x + 5) B. y = -1/2(x – 5) C. y – 5 = 0(x – 1/2) D. y – 5 = -1/2(x – 0)